OpenStudy (anonymous):

quarter-wavelength stubs of transmission lines are used as impedance- matching transformers. the formula for the proper characteristic impedance of matching the section is Zmatching= sqrt{Zlines x Zload}, where Zline is the characteristic impendance of the transmission line and Zload is the impedance of the load which must be matched to the transmission line. Assume that a transmission line having a characteristic impendance 300ohms is matched to an unknown load. You know that the characteristic impedance of a quarter-wave matching section is 72 ohms. What is the impendance of the load? 1) 125

OpenStudy (anonymous):

1)1250ohms 2)172.8ohms 3)35.36ohms 4)353.6ohms 5)125.0ohms 6)17.28ohms

OpenStudy (anonymous):

Zlines=300ohms Zload=? Zmatching=72ohms

OpenStudy (anonymous):

\[72=\sqrt{300 x}\]

OpenStudy (anonymous):

\[72^2=300 x\]

OpenStudy (anonymous):

5184 = 300 x x=432/25

OpenStudy (anonymous):

17.28

OpenStudy (anonymous):

how did you get the 432/25?

OpenStudy (anonymous):

5184 = 300 x \[\frac{5184}{300}=x\]

OpenStudy (anonymous):

hum i still dont really understand you divided 5184 by 300 and that comes to 17.28 but how did you get 432 divided but 25?

OpenStudy (anonymous):

simplifying \[\frac{5184}{300}\] into \[\frac{432}{25}\]